\newpage
\begin{thebibliography}{99}

\bibitem{Borwein}
The Riemann Hypothesis: \textbf{A Resource for the Afficinado and
Virtuoso Alike}. Borwein, Choi, Rooney and Weirathmueller
(Editors), Springer (2008); \\
The theory of the Riemann Zeta-Function, Second Edition, E.C.
Titchmarsh, revised by D. R. Heath-Brown, Clarendon Press, Oxford
(1986).

\bibitem{Barford}
T. Barford, arXiv: nucl-th/0404072 v 1, 26 April 2004.

\bibitem{Khuri}
N. N. Khuri: Inverse Scattering, the Coupling Constant Spectrum
and the Riemann Hypothesis, arXiv: hep-th/0111067 v 1, 7 November
2001; \\
K. Chadan and P. C. Sabatier: Inverse Problems in Quantum Scattering
Theory, Second Edition, Springer-Verlag (1989); \\
R. G. Newton: Inverse Schr\"{o}dinger Scattering in 3 dimensions,
Springer-Verlag (1989).

\bibitem{Krutov}
A. F. Krutov, D. I. Muravyev, and V. E. Troitsky:
A Factorization of the $S$-matrix into Jost Matrices,
\emph{J. Math. Phys.} \textbf{38} (1997), 2880.

\bibitem{Levin}
B. J. Levin: Distribution of Zeros of Entire Functions,
American Math. Soc. Traslations of Math monographs, Vol. 5,
Providence, RI, 1964. \\
See also, B. Ghusayni, Entire Functions of Order One and Infinite
Type (preprint).

\bibitem{Okubo}
S. Okubo: Lorentz Invariant Hamiltonian and Riemann Hypothesis,
\emph{J. Phys A: Math Gen} \textbf{31} (1998), 1049.

\bibitem{Mezzadri}
Recent Perspectives in Random Matrix Theory and Number Theory. F.
Mezzadri and N.C. Snaith (Editors), Cambridge University Press
(2005);\\
G. Sierra: A Quantum Mechanical Model of the Riemann Zeros, arXiv:
math-ph/07111063;\\
M. V. Berry and J. P. Keating: H=xp and the Riemann Zeros (1999); \\
Jean-Francois Burnol, arXiv: math.NT 08010530 v 2, 4 January 2008.

\bibitem{Castro}
Carlos Castro: \emph{Int. Jour. Geom. Methods in Modern Physics}
Vol. \textbf{5} (2008), 17. \\
H. C. Rosu: \emph{Mod. Phys. Lett.} A \textbf{18} (2003), 1205

\bibitem{Gradshteyn}
I. S. Gradshteyn and I. M. Ryzhik: Table of Integrals, Series, and
Products, Corrected and Enlarged Edition, Academic Press (1980),
page 693.

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{amsmath, amssymb}

\begin{document}

\title{
REALIZATION OF THE \\
RIEMANN HYPOTHESIS \\
VIA \\
COUPLING CONSTANT SPECTRUM}

\author{R. Acharya \footnote{I dedicate this note to my teacher, George Sudarshan
of the University of Texas at Austin}\\
Department of Physics\\
Arizona State University\\
Tempe, AZ 85287-1504}

\maketitle
\thispagestyle{empty}

\begin{abstract}
We present a Non-relativistic Quantum mechanical model, which exhibits the
realization of Riemann Conjecture. The technique depends on
exposing the $S$-wave Jost function at zero energy and in
identifying it with the Riemann $\xi(s)$ function following a
seminal paper of N. N. Khuri.
\end{abstract}

\newpage
\setcounter{page}{1}

We begin by recalling the all-too-familiar lore that the Riemann
hypothesis has been the Holy Grail of mathematics and physics for
more than a century \cite{Borwein}. It asserts that \textbf{all} the zeros of
$\xi(s)$ have $\sigma = \frac{1}{2}$, where $s=\sigma \pm it_{n}$,
$n = 1, 2, 3 \dots \infty$. It is believed all zeros of $\xi(s)$
are simple. The function $\zeta(s)$ is related to the Riemann
$\xi(s)$ function via the defining relation \cite{Borwein},

\begin{equation}\label{eq:1}
\xi(s) = \frac{1}{2} s(s-1)\pi^{-\frac{S}{2}} \Gamma \Bigl( \frac{s}{2} \Bigr)
\zeta(s)
\end{equation}
so that $\xi(s)$ is an entire function, where

\begin{equation}\label{eq:2}
\zeta(s) = \sum^{\infty}_{n=1} n^{-s}, \quad s = \sigma + it, \quad \sigma >
1
\end{equation}

$\zeta(s)$ is holomorphic for $\sigma > 1$ and can have No zeros for
$\sigma > 1$. Since $1/\Gamma(z)$ is entire, the function $\Gamma
\bigl( \frac{s}{2} \bigr)$ is non-vanishing, it is clear that
$\xi(s)$ \textbf{also} has no zeros in $\sigma > 1$: the zeros
of $\xi(s)$ are confined to the ``critical strip" $0 \leqslant \sigma
\leqslant 1$. Moreover, if $\rho$ is a zero of $\xi(s)$, then so is
$1-\rho$ and since $\overline{\xi(s)} = \xi(\overline{s})$, one deduces that
$\overline{\rho}$ and $1 - \overline{\rho}$ are also zeros. Thus the Riemann
zeros are symmetrically arranged about the real axis
\textbf{and} also about the ``critical line" given by $\sigma =
\frac{1}{2}$. The Riemann Hypothesis, then, asserts that ALL zeros
of $\xi(s)$ have Re $s= \sigma = \frac{1}{2}$.

We conclude this introductory, well-known remarks with the
assertion that every entire function $f(z)$ of \textbf{order
one} and ``\textbf{infinite type}" (which guarantees the
existence of \textbf{infinitely many Non-zero}
zeros can be represented by the Hadamard factorization, to wit \cite{Borwein},

\begin{equation}\label{eq:3}
f(z) = z^{m} e^{A} e^{BZ} \prod^{\infty}_{n=1}
\Bigl( 1-\frac{z}{z_{n}} \Bigr) \exp \Bigl( \frac{z}{z_{n}} \Bigr)
\end{equation}
where `m' is the multiplicity of the zeros (so that $m=0$, for
simple zero).

Finally, $\xi(s) = \xi(1-s)$ is indeed an entire function of order
one and infinite type and it has \textbf{No zeros either for
$\mathbf{\sigma > 1}$ or $\mathbf{\sigma < 0}$.}

We begin our brief note by defining the non-relativistic, quantum
mechanical potential model in 3 dimensions.

The zero energy (i.e., $k^{2} = 0$, $\frac{2m}{\hbar^{2}} = 1$)
Schrodinger equation reads [for \textbf{zero} angular momentum,
$l = 0$ ($S$-waves)]

\begin{equation}\label{eq:4}
\frac{1}{r^{2}} \frac{\partial}{\partial r} \biggl( r^{2}
\frac{\partial \psi(r)}{\partial r} \biggr) + \biggl(
0 - \frac{\lambda}{r^2} \biggr) \psi(r) = 0
\end{equation}

\textbf{The potential we choose is:}
\begin{equation*}
\tag{\ref{eq:4}$'$}
\begin{split}
V(r) = \frac{\lambda}{r^{2}}
\end{split}
\end{equation*}

The canonical change of the independent variable $\psi(r)$ to
$\frac{U(r)}{r}$ results in the equation, well-known to every one
(!):

\begin{equation}\label{eq:5}
U''(r) - \frac{\lambda}{r^{2}} U(r) = 0
\end{equation}

Eq. (\ref{eq:5}) has the solution:

\begin{equation}\label{eq:6}
U(r) = r^{s} \Rightarrow \psi(r) = r^{s-1}
\end{equation}
where $s$ is \textbf{constrained} to satisfy the relation,

\begin{equation}\label{eq:7}
\lambda = s (s-1)
\end{equation}

For the Regular solution at the origin, i.e., $\psi(0) = 0$, we
require that Real $s > 1$. Eq. (\ref{eq:7}) will play a crucial role in the
upcoming analysis. We note in passing that the Repulsive nature of
the real potential, i.e., $V(r) = \frac{\lambda}{r^{2}}$,
$\lambda$ real and $> 0$, requires that there are \textbf{No
bound states.} Further due to the scale invariance of the problem,
i.e., the potential is scale-free (and we are suppressing the
overall factor $\frac{1}{M}$ in $V(r)$). The \textbf{coupling
constant $\mathbf{\lambda}$ is dimensionless.}

The inverse-square potential acts like a centrifugal term in the
free Schro-dinger equation. It is a straightforward exercise in
undergraduate physics to determine the corresponding $S$-wave
phase shift for scattering solution \cite{Barford}:

\begin{equation}\label{eq:8}
\delta = \frac{\pi}{4} - \frac{\pi\nu}{2}
\end{equation}
where

\begin{equation}\label{eq:9}
\nu = \sqrt{\frac{1}{4} + \lambda} > \frac{1}{2}
\end{equation}

We note in passing that for the free case, $\lambda = 0$, $\nu =
\frac{1}{2}$ and the phase shift does indeed vanish and the $S$
matrix is unity:
\begin{equation}\label{eq:10}
S = \exp(2i\delta)
\end{equation}

It is important to emphasize that the phase shift $\delta$ in Eq.
(\ref{eq:8}) and hence also the $l = 0$ partial wave scattering
matrix in Eq. (\ref{eq:10}) are \textbf{independent of energy} and
the $S$ matrix is thus manifestly scale invariant. (There are no
bound states).

We go on to summarize the seminal idea of Khuri (and Chadan)
\cite{Khuri}. It is well known that the $S$ wave Jost function is
identified via the defining relation,

\begin{equation}\label{eq:11}
S(k) = \frac{F_{-}(k)}{F_{+}(k)}
\end{equation}
where $F_{+}(k)$ is holomorphic in the upper half complex $k$
plane and $F_{-}(k)$ is holomorphic in the lower half of $k$
plane, where $k$ is the momentum (in the center-of-mass frame).

Thus,

\begin{equation}\label{eq:12}
\begin{split}
F_{+}(k) = {S}^{-1} (k) F_{-}(k), \qquad  & Im k = 0 \\
                                         & -\infty < k < \infty
\end{split}
\end{equation}
and

\begin{equation}\label{eq:13}
\det S(k) \neq 0, \qquad Imk = 0
\end{equation}

Khuri's idea is the following: the $S$-wave Jost function for a
potential

\begin{equation}\label{eq:14}
\lambda V^{\ast} = \frac{\lambda}{r^{2}}
\end{equation}
is an entire function of $\lambda$ with an infinite number of
zeros extending to infinity. For a repulsive potential $V$ and at
\textbf{zero energy}, these zeros of the coupling constant
$\lambda$, will all be \textbf{real} and \textbf{negative}
(see the elaboration of this fact later on), i.e.,
$\lambda_{n}(k^{2} = 0) < 0$.

By rescaling $\lambda$ such that $\lambda_{n} < -\frac{1}{4}$ and
changing variables to $s$, with $\lambda = s(s-1)$, it follows
that as a function of $s$, the $S$-wave Jost function has
\textbf{only} (!) zeros on the line $s_{n} = \frac{1}{2} +
it_{n}$. Thus, if we can ``find" a \textbf{repulsive} potential
$V^{\ast}$ whose coupling constant spectrum coincides with the Riemann
zeros of $\xi(s)$, this will unambiguously establish the holy
grail of physics and mathematics! Khuri commented that
\textbf{``This will be a very difficult and unguided search."}
That said, we wish to make the crucial observation: For the
potential at hand, i.e., $V(r) = \lambda V^{\ast}(r)$,
$V^{\ast}(r)=\frac{1}{r^{2}}$, $\lambda$ real \& positive, the
coupling constant $\lambda$ must \textbf{necessarily} satisfy
the equality,

\begin{equation}\label{eq:15}
\lambda = s(s-1)
\end{equation}

In other words, there is ``no need to change variables!" We
proceed by summarizing the ``solution" to the boundary value
problem defined by Eq. (\ref{eq:12}) for the $S$ wave Jost function.

In the case of uncoupled partial waves the solution to the
boundary value problem formulated via Eq. (\ref{eq:12}) has been
given by Krutov, Muravyev and Troitsky \cite{Krutov} in 1996. The
``Solution" reads:

\begin{equation}\label{eq:16}
F_{\pm}(k) = \Pi_{\pm}(k) \exp \biggl(
\frac{1}{2\pi i} \int^{\infty}_{-\infty}
\frac{\ln(S^{-1}(k) \Pi^{2}_{-}(k))}{k'-k\mp io}dk'
\biggr)
\end{equation}
where

\begin{equation}\label{eq:17}
\Pi_{\pm}(k) = \prod^{m'}_{j=1} \frac{k \mp ik_{j}}{k \pm ik_{j}},
\quad k_{j} >0, \quad j=1,2,\dots,m' \quad (m' < \infty)
\end{equation}

\begin{equation}\label{eq:18}
\Pi_{\pm}(k) \equiv 1, \quad m' = 0
\end{equation}
where $m'$ is the number of bound states.

It is well-known, of course that

\begin{equation}\label{eq:19}
\frac{1}{k'-k \mp io} = P \frac{1}{k'-k} \pm i\pi \delta(k'-k)
\end{equation}

Proceeding further, we set the Jost function ($S$ wave), at zero
energy

\begin{equation}\label{eq:20}
F_{+}(s(s-1); k^{2} = 0 ) \equiv \chi(s)
\end{equation}

The simplification of Eq. (\ref{eq:16}) is immediate. One finds
easily that

\begin{equation}\label{eq:21}
F_{+} = \chi(s) = \exp(-2i\delta), \quad S(s) = \exp(2i\delta)
\quad \text{and} \quad F_{-}(s) = 1
\end{equation}

We note in passing that since the partial wave $S$ matrix is
independent of $k$ (energy) due to scale invariance, one can set
$k=0$ in simplifying Eq. (\ref{eq:16}). Recall that there are no
bound states either, $m' = 0$ and $\Pi_{\pm}(k) \equiv 1$.

Since $\chi(s)$ defined in Eq. (\ref{eq:20}) is entire \cite{Levin}
and has \textbf{only zeros} on the real line Re $s_{n} =
\frac{1}{2}$ (see clarification of this later) and the number of
non-zero zeros are infinite on the ``critical" line Re $s_{n} =
\frac{1}{2}$, the Hadamard factorization Eq. (\ref{eq:3}) is valid,
i.e.,

\begin{equation}\label{eq:22}
\chi(s) = e^{A} e^{Bs} \prod^{\infty}_{n=1} \biggl( 1-\frac{s}{s_{n}}
\biggr)\exp\biggl(\frac{s}{s_{n}}\biggr)
\end{equation}

It will turn out that both the constants $A$ and $B$ can be
identified (see below) and also that $\chi(s)$ has No zeros either
for $s > 1$ or $s < 0$.

From Eq. (\ref{eq:8}), Eq. (\ref{eq:9}) and Eq. (\ref{eq:10}), we
obtain

\begin{equation}\label{eq:23}
\nu = \frac{1}{2} - \frac{2}{\pi} \delta = \frac{1}{2} +
\frac{1}{\pi} |\ln\chi(s)| > \frac{1}{2}
\end{equation}
where we have set (Eq. (\ref{eq:21}))

\begin{equation}\label{eq:24}
\ln \chi(s) = +i|\ln \chi(s)|
\end{equation}
in order to satisfy that $\delta < 0$

\begin{equation}\label{eq:25}
\delta = - \frac{1}{2} |\ln \chi(s)|
\end{equation}

[Recall that for a \textbf{repulsive} potential, $\lambda > 0$ the
phase shift has to be \textbf{negative}!]

\begin{equation}\label{eq:26}
\therefore \lambda = s(s-1) = \nu^{2} - \frac{1}{4}
\end{equation}
giving [From Eq. (\ref{eq:23}), Eq. (\ref{eq:25})]

\begin{equation}\label{eq:27}
\lambda = \frac{1}{\pi} | \ln \chi(s) | + \frac{1}{\pi^{2}} |\ln
\chi(s)|^{2} > 0
\end{equation}
where $\lambda$ is real positive and $s$ real and $> 1$.

From Eq. (\ref{eq:26}) and Eq. (\ref{eq:27}), we obtain $\mathbf{\chi(s)}$
\textbf{for real $\mathbf{s > 1}$:}

\begin{equation}\label{eq:28}
|\ln \chi(s)| = \pi(s-1) > 0 \qquad s \text{, real \&} > 1
\end{equation}

\begin{align}
\therefore \chi(s) & = e^{\pi(s-1)}, \quad \mathbf{s > 1} \textbf{,
real} \label{eq:29} \\
                   & = e^{-\pi s}, \qquad \mathbf{s < 0} \textbf{,
                   real} \label{eq:30}
\end{align}

$A = - \pi$, $B = \pi$ for $s > 1$, real (from Eq. (\ref{eq:22}))

Notice, as promised that $\chi(s)$ is entire (and order one) and
\textbf{has No zeros for $\mathbf{s > 1}$ and $\mathbf{s < 0}$.}
It is worth noticing that since $\lambda = s(s-1)$ is invariant
under $s \rightarrow 1-s$, the forms of $\chi(s)$ in Eq.
(\ref{eq:29}) and Eq. ({\ref{eq:30}) are consistent.

We now comment on the \textbf{all-important} restriction that for
a real, repulsive potential the $S$-wave Jost function $\lambda V$
is an entire function of $\lambda$ with an infinite number of
zeros extending to infinity and these zeros will all be
\textbf{real and negative}. [This can be verified following Khuri
{See his Eq. (1.1) and Eq. (1.2)] i.e.,

\begin{equation}\label{eq:31}
[I m \lambda_{n} (i \tau)] \int^{\infty}_{0} | f (\lambda_{n}(i \tau));
i \tau; r |^{2} dr = 0
\end{equation}
where $k = i \tau$.

We observe that Eq. (\ref{eq:31}) does \textbf{not} have the
potential $V(r)$ under the integration sign. This comes about by
examining the Schrodinger equation for $f(k,r)$ [which is
identical to the case of one dimension] and its complex conjugate
$f^{*}(k,r)$, and by performing the ``canonical" operation, i.e.,
multiplying the equation for $f(k,r)$ by $f^{*}(k,r)$,
subtracting and then importantly multiplying the resulting
equation by $r^{2}$. This ``gets rid of" the potential term on
R.H.S. Subsequently, one integrates by parts the equation thus
obtained, making use of the fact that the Wronskian of $f(k,r)$
and $f^{*}(k,r)$ is \textbf{independent} of $r$ (because they
satisfy the \textbf{same} Schrodinger equation: See Chaden and
Sabatier, \cite{Khuri}). The end result is Eq. (\ref{eq:31})
above, the potential $V(r)$ \textbf{not} present (in contrast to
Khuri's Eq. (1.2)). The rationale behind this is to ensure that
the integral in Eq. (\ref{eq:31}) is \textbf{now finite}, under
the stated conditions (See Eq. (36$'$) and Eq. (36$''$). If the
potential $V(r) = \lambda \frac{1}{r^{2}}$ were present, then the
corresponding integral in Eq. (\ref{eq:31}) would diverge!

The crucial point to verify is that the integral in Eq.
(\ref{eq:31}) is \textbf{non vanishing} and \textbf{finite. Only
then,} can one conclude that

\begin{equation}\label{eq:32}
I m \lambda_{n} (i \tau) = 0.
\end{equation}

We proceed to demonstrate this as follows.

The $S$ wave Jost \textbf{solution} $f(k,r)$ for the
potential $V(r) = \frac{\lambda}{r^{2}}$ ($\lambda >
0$) is well-known \cite{Barford}:

\begin{equation}\label{eq:33}
f(k,r) = \sqrt{\frac{\pi k r}{2}} e^{i(\frac{\pi}{2}\nu +
\frac{\pi}{4})} H^{(1)}_{\nu} (kr)
\end{equation}
where $H^{(1)}_{\nu} (kr)$ is Hankel function of $I$ kind.

One \textbf{cannot} obtain the Jost function by taking the limit
$r \rightarrow 0$ in Eq. (\ref{eq:33}) because this limit does
\textbf{not} exist. This only works for REGULAR POTENTIALS! (See
Chadan and Sabatier, Page 10, Eq. (I.3.6).)
We can bypass this conundrum, because we
already ``know" the $S$ wave Jost function $F_{+}(k) = F_{+}(0)$
from Eqs. (\ref{eq:8}), (\ref{eq:9}) and (\ref{eq:21}).

We observe in passing that the Jost solution Eq. (\ref{eq:33}) has
the required asymptotic behavior, i.e.,

\begin{equation}\label{eq:34}
f(k,r) \mathrel{\mathop{\longrightarrow}\limits_{r\rightarrow \infty}} e^{ikr}
\end{equation}

Since
\begin{equation}\label{eq:35}
H^{(1)}_{\nu}(z) \sim \sqrt{\frac{2}{\pi z}}
e^{i(z-\frac{\pi}{2} \nu - \frac{\pi}{4})},
\end{equation}
it is now straightforward to check that
\begin{equation}\label{eq:36}
\int^{\infty}_{0} |f(\lambda_{n}(i\tau); i\tau; r|^{2} dr < \infty
\end{equation}
by plugging in Eq. (\ref{eq:33}) for $f(k,r)$.

\begin{equation*}
\tag*{Eq. (4$'$)}
\begin{split}
V(r) = \frac{\lambda}{r^{2}}
\end{split}
\end{equation*}

Eq. (\ref{eq:31}), (\ref{eq:33}) give

\begin{equation*}
[ I m \lambda_{n} (i\tau)\tau ] \frac{2}{\pi}
\int^{\infty}_{0} r K^{2}_{\nu} (\tau r) dr = 0
\tag{\ref{eq:36}$'$}
\end{equation*}
where
\begin{equation*}
\int^{\infty}_{0} r K^{2}_{\nu} (\tau r) dr = \frac{1}{8}
\frac{1}{\tau^{2}} \frac{\pi \nu}{\sin \pi \nu}
\tag{\ref{eq:36}$''$}
\end{equation*}

$\mathbf{\therefore}$ \textbf{Eq. (\ref{eq:32}) follows from Eq. (36$'$), for $\mathbf{\tau
\neq 0}$ ($\mathbf{\tau > 0}$) \cite{Gradshteyn}}

\begin{equation*}
I m \lambda_{n} (i \tau) = 0, \qquad \nu > \frac{1}{2}, \quad \nu \neq 1, 2, 3 \dots \infty. \tag{32}
\end{equation*}

We now follow Khuri \cite{Khuri}:

All zeros of $\lambda_{n}(i\tau)$ must \textbf{be real} ($\tau >
0$) ``But $\lambda_{n}(i\tau)$ must be negative since the
potential $[\lambda_{n}(i \tau) V^{\ast}]$ will have a bound state
at $E = - \tau^{2}$ and that could not happen if $V^{\ast}
\geqslant 0$ and $\lambda_{n}(i \tau) > 0$. Hence by continuity,
$\mathbf{\lambda_{n}(0)}$\textbf{, for all $\mathbf{n}$, is real and
negative.} The zero energy coupling spectrum lies on the negative real
line for $V^{\ast} \geqslant 0$." In our case, in view of Eq.
(\ref{eq:7}), i.e.,
\begin{center}
$\lambda = s(s-1)$ is mandatory!
\end{center}

\textbf{Thus, we conclude unequivocally that $\mathbf{\chi(s)}$
has all its infinite number of nonzero zeros on the critical
line. And can be set equal to Riemann's $\mathbf{\xi}$
function!}

This establishes the Holy grail of physics and mathematics!!

We conclude by commenting on the precise relation between the $S$
wave Jost function, $\chi(s)$ and Riemann's $\xi(s)$ function.
Both are entire, order one and infinite type and possess identical
infinite number of non-zero zeros \textbf{only} on the critical
line Re $s = \frac{1}{2}$ and both have no zeros either for
$\sigma > 1$ or $\sigma < 0$. Their ratio, by Hadamard's theorem
is given by

\begin{equation}\label{eq:37}
\frac{\chi (s)}{\xi(s)} = e^{\alpha + \beta s}, \quad \text{all} \; s
\end{equation}

Since \cite{Mezzadri}

\begin{equation}\label{eq:38}
\xi (s) = \frac{1}{2} e^{b_{0}s} \prod_{n} \biggl(
1-\frac{s}{s_{n}} \biggr) e^{\frac{s}{s_{n}}}, \quad \textbf{all}
\; s
\end{equation}
where

\begin{equation}\label{eq:39}
b_{0} = - \frac{1}{2} \gamma - 1 - \frac{1}{2} \ln 4 \pi
\end{equation}
\begin{center}
($\gamma$ is Euler's constant)
\end{center}

We obtain the identification,

\begin{align}
          \alpha & = \ln2 - \pi \label{eq:40} \\
\text{and} \quad \beta & = \pi + \frac{1}{2} \gamma + 1 - \frac{1}{2}
\ln 4\pi \label{eq:41}
\end{align}

I wish to thank Professor Coffey, Professor Odlyzko and Professor
Ivic for a communication on the zeros of $\xi(s)$. I am especially
grateful to Professor Okubo \cite{Okubo} for informing about his
work on the Riemann Hypothesis. I am grateful for a communication
from Professor Carlos Castro and Professor H. C. Rosu. I am grateful to Irina Long for
her generous help in the preparation of this paper.

\end{document}